### 07 Feb 2024 v1

@@ 3.2 Implementation procedure
The overview is shown in Fig.5.The development process is divided into 4 sub-stages: 1. We utilize S2orc as our infrastructure data of KG. S2orc is an academic database that encompasses metadata, bibliographic references, and full text for 8.1 million open access papers[1]. 1. In the first stage, we perform secondary processing based on the S2ORC database to build our insight survey dataset. 2. In the second stage, we use the Sentence Bert model and manually set labels to perform three-class classification (Resolved/Neutral/Finding) on the insight content of each paper to extract the corresponding sentence.3. In the third stage, we use the extracted sentences from phase 2 to perform in-depth analysis on inheritance and relevance chain. We select appropriate papers from the whole insight survey dataset based on certain criteria to generate hierarchical trees.4. In the final stage, we use the pyvis[15] to visualize the tree-structured graph and provide some case studies to demonstrate our findings. The overview is shown in Fig.5.
@@ 3.4 Phase2 : Insight Sentence Extraction
Sentence Segmentation: Spacy is an open-source natural language processing library for Python that offers an API to access its machine learning trained methods and properties[11]. In this work, we utilize the spacy pre-trained model to implement rule-based sentence segmentation. Spacy provides a pre-trained English library called ”en_core_web_sm” which includes a default sentence segmenter [^1] https://www.tutorialspoint.com/perform-sentence-segmentation-using-python-spacy. Any complex segmentation patterns that failed were manually fixed.
Human Annotation: For the segmented sentences mentioned above, experts determine the viewpoint of each sentence based on its meaning. The viewpoint include ’Resolved’, ’Neutral’, or ’Finding’. Table 2 specific definitions, distinguishing criteria, and examples of these three labels.
@@ 3.5 Phase3-4 : Hierarchical Tree Construction & Visualization
To provide researchers with a visual understanding(Fig.1-Fig.2) of the guidance on research directions provided by these selected papers. We use pyvis to visualize KG based on the configuration of Table 1. In addition to the co-occurring vocabulary of insight sentences between papers, we also extract keywords specific to each paper to capture its unique concept. When selecting keywords for each paper in the KG, we combine the ‘insight content’ from all papers in the subtree to create the overall text. The ’insight content’ of a specific article within the subtree is used as the partial text. Tf-idf[13] method used to extract partial text keywords from the global text composed of multiple paper-texts in a tree, and then display them as paper keywords in the node titles, and a part of the representative sample branch included insight sentences is shown in Fig.5.
3.6 Case Study & Analysis
The detail of sample branches in Fig.1-Fig.2 are shown in Fig.5-Fig.5, we observe that the occurrence probability of co-occurring words in the ’insight sentences’ is relatively low, as indicated by the edges of the Inheritance tree. This is because the Inheritance tree is based on reference extraction, and the citations of previous studies are often related to the research background or methodology. This reflects the weakness of correlation in the insight survey. In contrast, the papers selected by the relevance tree exhibit higher occurrence probability and similarity of co-occurring words in the ’insight sentences’. Although the chain from ”Issue finding” to ”Issue Resolved” may not be completely reasonable in terms of content, it provides researchers with certain keywords and guidance for research directions. On the other hand, the inheritance tree serves as a broader extension based on a specific research topic. However, there is a need to incorporate the similarity of relevance between papers horizontally in order to expand the breadth of the KG and enhance the effectiveness of the inheritance tree in supporting researchers’ understanding of research theories and relevance.
Figure 4: Inheritance tree Sample branch(p2,p4,p5 in Fig.1)
Figure 4: Inheritance tree Sample branch(p2,p4,p5 in Fig.1)
Figure 5: Relevance tree Sample branch(p2,p4,p5 in Fig.2)

### 07 Apr 2024 v2

@@ 3.2 Implementation procedure
The overview is shown in Fig.3. The development process is divided into 4 sub-stages: 1. We utilize S2orc as our infrastructure data of KG. S2orc is an academic database that encompasses metadata, bibliographic references, and full text for 8.1 million open access papers[1]. 1. In the first stage, we perform secondary processing based on the S2ORC database to build our insight survey dataset. 2. In the second stage, we use the Sentence Bert model and manually set labels to perform three-class classification (Resolved/Neutral/Finding) on the insight content of each paper to extract the corresponding sentence.3. In the third stage, we use the extracted sentences from phase 2 to perform in-depth analysis on inheritance and relevance chain. We select appropriate papers from the whole insight survey dataset based on certain criteria to generate hierarchical trees.4. In the final stage, we use the pyvis[15] to visualize the tree-structured graph and provide some case studies to demonstrate our findings.
@@ 3.4 Phase2 : Insight Sentence Extraction
Sentence Segmentation: Spacy is an open-source natural language processing library for Python that offers an API to access its machine learning trained methods and properties[11]. This work uses the spacy pre-trained model to implement sentence segmentation. Spacy provides a pre-trained English library called en_core_sci_lg which includes a default sentence segmenter [^1] https://allenai.github.io/scispacy/ [^2] https://www.tutorialspoint.com/perform-sentence-segmentation-using-python-spacy. Any complex segmentation patterns that failed were manually fixed.
Human Annotation: For the segmented sentences mentioned above, experts determine the viewpoint of each sentence based on its meaning. The viewpoints include ’Issue Resolved,’ ’Neutral,’ or ’Issue Finding.’ Sentences that thoroughly analyze research methods without explicitly highlighting contributions will be considered ‘Neutral.’ Sentences discussing potential future works in a particular field without specificity will also fall into the ‘Neutral’ category. ’Issue Resolved‘ sentences need at least a combination of contribution, research details, and experimental outcomes, regardless of their positive or negative outcomes. Sentences that ambiguously hint at trends and recommendations may be classified as ‘Issue Finding.’ When categorizing these three viewpoints, the references, tables, and figures of the sentences are unchanged. Table 2 shows specific definitions, distinguishing criteria, and examples of these three labels. The dataset, which consists of insight sentences and labels, has been published on  [^3] https://www.kaggle.com/datasets/dannyleeakira/dataset-for-academic-novelty-insight-survey
@@ 3.5 Phase3-4 : Hierarchical Tree Construction & Visualization
To provide researchers with a visual understanding(Fig.1-Fig.2) of the guidance on research directions provided by these selected papers. We use pyvis to visualize KG based on the configuration of Table 1. In addition to the co-occurring vocabulary of insight sentences between papers, we also extract keywords specific to each paper to capture its unique concept. When selecting keywords for each paper in the KG, we combine the ‘insight content’ from all papers in the subtree to create the overall text. The ’insight content’ of a specific article within the subtree is used as the partial text. Tf-idf[13] method used to extract partial text keywords from the global text composed of multiple paper-texts in a tree, and then display them as paper keywords in the node titles, and a part of the representative sample branch included insight sentences is shown in Fig.LABEL:figure3.